15.4 Is HDTV ownership related to quantity of purchases of other electronics? A Best Buy retail outlet collected the followinPlease state your hypotheses, test statistic, pvalue and make a definitive conclusion statement.

Please show all your work for my understanding! Thank you so much
Answer #1

Ans:

H0: In-store purchases is independent of the number if large-screen HDTVs owned.

Ha: In-store purchases is dependent on the number if large-screen HDTVs owned.

Observed(fo)
None One more than one Total
None 13 14 13 40
One 18 31 31 80
Two or more 19 45 66 130
Total 50 90 110 250
Expected(fe)
None One more than one Total
None 8.0 14.4 17.6 40.0
One 16.0 28.8 35.2 80.0
Two or more 26.0 46.8 57.2 130.0
Total 50 90 110 250
Chi square=(fo-fe)^2/fe
None One more than one Total
None 3.125 0.011 1.202 4.338
One 0.250 0.168 0.501 0.919
Two or more 1.885 0.069 1.354 3.308
Total 5.260 0.248 3.057 8.565

Test statistic:

Chi square=8.565

df=(3-1)*(3-1)=4

p-value=chidist(8.565,4)=0.0729

As,p-value<0.10,we reject the null hypothesis.

There is sufficient evidence to conclude that In-store purchases is dependent on the number if large-screen HDTVs owned.